A farmer grows a kilogram of wheat and sells it to a Miller for $1.00. The Miller turns the 

wheat into flour and then sells the flour to a baker for $3.00. The baker uses the flour to 

make bread and sells the bread to an engineer for $6.00. The engineer eats the bread. What is 

the value added by each person? What is GDP?



1
Expert's answer
2021-07-06T18:25:22-0400

Each person adds:


The farmer: "1-0=1"


The miller: "3-1=2"


The baker: "6-3=3"


GDP= final product cost=6 or GDP= amount of value added="1+2+3=6"